# Learning Robust Intervention Representations with Delta Embeddings

### v1

@@ Introduction
    where $P(i)$ is the set of all positive samples for anchor $i$ in the batch, $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity, and $\tau$ is a scalar temperature hyperparameter. This loss component is also consistent with the structural equation (3).
  3. 3.
    Sparsity Regularizer: To encourage a minimal representation in line with the sparse mechanism shift hypothesis (Property 2, Sparsity), we apply an $\ell_{1}$ regularization penalty. This loss penalizes the sum of the absolute values of the embedding dimensions, promoting solutions where most dimensions are zero.
@@
where $\alpha_{\text{contrast}}$ and $\alpha_{\text{sparsity}}$ are scalar hyperparameters that balance the influence of each loss component.
Notice that although no loss component explicitly enforces Property 1, this property is directly satisfied by the use of the Delta Embedding and image pairs, where the observed changes are only due to $a$. This does not hold under the actionable counterfactual case (Liu et al. 2023), where additional scene variables may change across observations. If, however, these changes are not spuriously correlated with the action $a$ in the data, then the use of $\mathcal{L}_{CE}$ can still retrieve the CDE representation. A deeper investigation of the actionable counterfactual case is left as future work.
Spatial Extension: The Patch-Wise Model
Motivation
@@
  1. 1.
    Patch-wise Feature Extraction: We use a ViT backbone, but instead of taking the global ‘[CLS]’ token, we retain the output feature vectors for each individual image patch. This gives us a sequence of patch features for both the before and after images.
  2. 2.
    Patch-wise Delta Computation: A shared Causal Projector and the subtraction operation are applied independently to each corresponding pair of patch features. This yields a set of delta embeddings, $\{\delta_{p}\}$, one for each spatial patch location $p$.
  3. 3.
    The Aggregation Module ($\mathcal{G}$): Since the action classifier needs a single input vector, we must aggregate the set of patch-wise deltas. This module’s task is to identify the region of change and produce a single, representative delta vector $\bar{\delta}$. We employed the following simple and effective aggregation strategy:
    Top-K Aggregation: This strategy is based on the assumption that the action’s primary effect is localized to a few patches. We identify the $k$ patches with the largest change by measuring the $L_{2}$ norm of their delta vectors ($\|\delta_{p}\|_{2}$). The final delta $\bar{\delta}$ is the average of these top $k$ patch deltas.
The same loss function ($\mathcal{L}_{\text{total}}$) is then applied to the aggregated delta vector $\bar{\delta}$.
Experiments
@@
Experimental Setup
Our evaluation is conducted on the Causal Triplet benchmark (Liu et al. 2023), specifically designed for intervention-centric causal representation learning. This benchmark features three distinct settings of increasing visual complexity: single-object synthetic scenes, multi-object synthetic scenes (both from ProcTHOR (Deitke et al. 2022)), and challenging real-world scenes from Epic-Kitchens (Damen et al. 2022). In all settings models are trained on pairs of pre- and post-intervention images with action labels and are evaluated for their ability to infer the action. Further details on the datasets and data filtering procedures are provided in the Appendix.
@@
We compare our two proposed models against the baselines from the Causal Triplet paper (Liu et al. 2023), including vanilla ResNets (He et al. 2016), methods incorporating causal regularization (ICM, SMS), and object-centric approaches (Slot Attention (Locatello et al. 2020b), GroupViT (Xu et al. 2022)).
Main Quantitative Results
Table 1: Single-object ProcTHOR results. Our Global Delta Embedding model significantly improves OOD generalization under both compositional and systematic shifts. (R: ResNet-18, V: Vit-Small)
@@
Our CDE framework consistently delivers substantial improvements in OOD robustness across all evaluation settings, establishing a new state of the art. Our models show remarkable OOD generalization across all evaluation settings. For single-object scenes, our global CDE model cuts the generalization gap from 0.56 to 0.18 while matching IID accuracy (Table 1). In challenging multi-object and real-world settings (Tables 2 and 3), our Patch-Wise model outperforms all baselines, including oracle methods that use ground-truth segmentations masks. This validates our unsupervised approach, which boosts OOD accuracy by 6 points on the noisy Epic-Kitchens dataset.
Table 2: Multi-object ProcTHOR results (systematic shift). Our Patch-Wise model outperforms all baselines, including oracle methods.
Method IID Acc. OOD Acc. Gap ($\downarrow$)
ResNet $0.83_{\pm 0.01}$ $0.30_{\pm 0.08}$ $0.53$
@@
Ours(Patch-wise) 0.92±0.01 0.45±0.02 $0.47$
In summary, across all three levels of complexity, our Causal Delta Embedding framework consistently delivers substantial improvements in OOD robustness, showcasing its ability to learn invariant representations of causal actions.
Table 3: Real-world Epic-Kitchens results (systematic shift). Our method shows robust performance improvements in a challenging, unconstrained setting.
Method IID Acc. OOD Acc. Gap ($\downarrow$)
ResNet $0.42_{\pm 0.03}$ $0.17_{\pm 0.03}$ $0.25$
@@
Further Analysis
While our model demonstrates state-of-the-art out-of-distribution (OOD) accuracy, a deeper analysis is required to verify that this performance stems from the successful implementation of the theoretical properties of Causal Delta Embeddings.
Action Relationships in Causal Delta Space
To study the semantic structure of the learned delta space, we investigated whether the model could discover fundamental relationships between actions on its own. We computed the pairwise cosine similarity between all learned action representations and visualized the result in a similarity matrix (Figure 4).
The analysis reveals the following property: the model has learned a perfect anti-parallel relationship for opposite actions. The cosine similarity between the prototypes for open and close, for dirty and clean, and for turn on and turn off, is -1.0. This demonstrates that our framework not only separates the action concepts but also discovers opposing relationships between them, organizing the representations in a meaningful way. Furthermore we can see that a similar pattern occurs in the most challenging real-world dataset where the model learns the anti-parallel relationships for the actions open and close and also for fold and stretch (Figure 5).
Table 4: Ablation study of our method’s components on the ViT-Small model. Results are for the single-object systematic shift setting, showing the impact on OOD accuracy when each core component is removed.
@@
   w/o Sparsity Loss $0.96$ $0.70$
   w/o Contrastive Loss $0.95$ $0.60$
Baseline (CE Loss only) $0.94$ $0.59$
In summary, the combination of strong predictive properties and consistent semantic structure demonstrates that our CDE framework learns meaningful representations of interventions. For further geometric analysis of the delta space, including UMAP projections and k-NN classifier performance, refer to the Appendix.
Ablation Study
@@
To understand the contribution of each component of our CDE framework, we conducted a series of ablation studies. We analyze the impact of each major loss component on the performance of our primary model with a ViT-Small backbone. Table 4 presents the results, comparing our full model against versions where each theoretical component is removed, and a baseline trained only with standard CE loss.
@@
The results demonstrate the effectiveness of our approach. Our full model achieves an OOD accuracy of 73.073.0%, a +14+14 point improvement over the baseline trained solely with a CE objective, validating that explicitly structuring the representation space is critical for generalization. The supervised contrastive loss causes a 13-point drop in OOD accuracy when removed. The sparsity loss term causes a 3-point drop when removed.
For more ablation studies, refer to the Appendix.
Conclusion
@@
This paper introduces the Causal Delta Embedding (CDE) framework, a simple yet effective approach to interventional causal representation learning. By explicitly modeling interventions as delta vectors in a structured latent space, CDE inherently satisfies the properties of independence, sparsity and invariance, leading to improved generalization. Our empirical validation on the Causal Triplet challenge demonstrates that CDE achieves state-of-the-art OOD generalization, outperforming prior methods across synthetic and real world datasets. Beyond quantitative gains, we show that CDE learns semantically meaningful representations without supervision, where opposing actions have anti-parallel representations. Despite the promising results, we acknowledge that limitations remain for real-world data, since both IID and OOD accuracies are still low for real world deployment, and also the use of universal delta embeddings for each action limits its ability to capture context-dependent visual transformations of actions. Future research directions include dynamically identifying modified regions of the input states through attention mechanisms, extending the framework to video streams for modeling causal dynamics in temporal sequences, and investigating compositional properties of delta embeddings to enable multi-step interventions and generalization to novel action sequences.
@@ Appendix A Dataset Details
The ProcTHOR dataset (Deitke et al. 2022) provides synthetic indoor scenes. For our single-object scenes, each scene contains one manipulated object, ensuring a clear focus on the intervention. In multi-object scenes, multiple objects are present, increasing the visual complexity of the scene, although only one object is again manipulated. We follow the dataset generation and filtering procedures as described in (Liu et al. 2023) to ensure consistency with the Causal Triplet benchmark.
@@
The Epic-Kitchens dataset (Damen et al. 2022) comprises real world egocentric videos of diverse kitchen activities. From this, we extract pre- and post-intervention image pairs. Unlike synthetic environments, Epic-Kitchens introduces significant real world challenges such as camera motion, varying lighting conditions, occlusions and dynamic backgrounds, making the task of isolating interventions particularly challenging. To ensure dataset quality, a two-stage filtering process using Grounding DINO (Liu et al. 2024) for zero-shot object detection is applied. For each extracted pair, the pipeline verifies that the target object appears clearly in both frames with a detection confidence above a set threshold $t=0.45$. This automated filtering removes cases with poor object visibility or excessive motion blur.
Visualizations of OOD Shifts
Figures 6, 7 and 8 visually illustrate the compositional and systematic distribution shifts utilized in the Causal Triplet benchmark.
@@ Appendix B Geometric Analysis of Causal Delta Embeddings
Analysis of Learned Action Representations
@@
Figures 9 and 11 present the UMAP projection of individual delta embeddings from the IID and OOD test set of the single-object environment respectively. The delta embeddings in the IID setting achieve a clear separation between each action, leading to a near perfect IID accuracy as was presented by our experiments. On the other hand, while strong intra-class cohesion is visible, the global separation between these action clusters is not always visually distinct in the OOD setting. This suggests that while representations remain locally coherent, action representations are not as clearly discriminated compared to the IID setting. It is worth mentioning, however, that the 2D projection may not fully capture the features of the high-dimensional latent space.
k-NN Classifier Performance
Figure 10: Comparison of classifier performance on OOD test sets. The k-NN classifier, is able to classify OOD samples if trained with Delta Embeddings.
To quantitatively assess the quality of the local structure of the learned representations, we evaluate the performance of a simple, non-parametric k-Nearest Neighbors (k-NN) classifier directly on the Causal Delta Embeddings, where we set the number of neighbors $k=5$. A high k-NN accuracy indicates that the local neighborhoods are semantically meaningful and highly predictive of the action class. Figure 10 compares the k-NN classifier accuracy with the use of CDE with an MLP head (as in Figure 3) across all three benchmark settings. Although k-NN does not match the effectiveness of the MLP head, it still achieves comparable results in novel OOD samples, especially in the synthetic dataset. This provides an indication that the invariance property holds in the OOD case.
@@ Appendix C Ablation Studies
All the subsequent experiments ran on the single-object systematic shifts setting, except for the Top-K ablation study which ran on the multi-object systematic shifts setting.
@@
ViT-Small Ours (CE Only) $0.59$
ViT-Small Ours (Full Model) 0.73
@@
Impact of Backbone Architecture
To isolate the contribution of our CDE framework from the choice of feature extractor, we conducted a controlled comparison between our ViT-Small backbone and the ResNet-18 backbone used in the original Causal Triplet benchmark.
Impact of Loss Hyperparameters
@@ Appendix D Experimental Details
We do not apply any augmentations to the images, since we do not want to modify the interventional nature of the pairs. Augmentation in this problem could harm our assumptions. For example, a rotation could affect Equation (1) and eliminate the faithfulness of the encoder. We leave it as future work to investigate whether augmentations can boost OOD performance under different assumptions. We only resize images to $224\times 224$ pixels and apply zero-mean normalization with unit variance.
Optimization
@@
We use a batch size of 128 and an AdamW (Loshchilov and Hutter 2017) optimizer with a cosine annealing learning scheduler for 50 epochs. In the real world setting, we instead train for 100 epochs. The ViT feature extractor is not frozen but fine-tuned with a reduced learning rate of $10\%$ of the network’s base learning rate, which is set to $1\times 10^{-4}$. The weight decay parameter is $0.05$. All reported results include standard deviations computed over three independent runs with different random seeds.

### v2

@@ 5.1.2 Implementation of the Learning Objective
where $P(i)$ is the set of all positive samples for anchor $i$ in the batch, $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity, and $\tau$ is a scalar temperature hyperparameter. This loss component is also consistent with the structural equation 3. Finally, we introduce a (iii) Sparsity Regularizer: To encourage a minimal representation in line with the sparse mechanism shift hypothesis (Property 2, Sparsity), we apply an $\ell_{1}$ regularization penalty. This loss penalizes the sum of the absolute values of the embedding dimensions, promoting solutions where most dimensions are zero.
@@
where $\alpha_{\text{contrast}}$ and $\alpha_{\text{sparsity}}$ are scalar hyperparameters that balance the influence of each loss component. Note that the network is trained end-to-end, i.e., the encoder, $\phi$, is also updated during training.
5.1.3 Actionable counterfactual case
Notice that although no loss component explicitly enforces Property 1, this property is directly satisfied by the use of the Delta Embeddings and image pairs, where the observed changes are only due to $a$. This does not hold under the actionable counterfactual case (Liu et al., 2023), where the exogenous noise may change across observations, leading to $\epsilon\neq 0$ in equation 3 and non-zero elements in the delta embedding of equation 1. In the Appendix (Section E) we show that if $\epsilon$ is zero mean independent noise, then the classifier trained with $\mathcal{L}_{CE}$ is not affected by these non-zero values. This observation is also supported by our empirical findings presented in Section 6, where the delta embeddings lead to effectiveness improvements in real-world OOD data.
5.2 Spatial Extension: The Patch-Wise Model
@@ 5.2.1 Architecture
The architecture includes (i) Patch-wise Feature Extraction: We use a ViT backbone, but instead of taking the global ‘[CLS]’ token, we retain the output feature vectors for each individual image patch. This gives us a sequence of patch features for both the before and after images, (ii) Patch-wise Delta Computation: A shared Causal Projector and the subtraction operation are applied independently to each corresponding pair of patch features. This yields a set of delta embeddings, $\{\delta_{p}\}$, one for each spatial patch location $p$. (iii) Top-K selection: We assume that the action’s primary effect is localized to a small number of patches. We therefore identify the $k$ patches with the largest change by measuring the $L_{2}$ norm of their delta vectors ($\|\delta_{p}\|_{2}$). The loss $\mathcal{L}_{\text{total}}$ of equation 6 is then applied to each of the $k$ delta vectors, leading to one loss $\mathcal{L}^{(i)}_{\text{total}}$ for each vector. The final loss is then the average of the individual patch losses, $\mathcal{L}=\frac{1}{k}\sum_{i=1}^{k}\mathcal{L}^{(i)}_{\text{total}}$.
6 Experiments
@@
6.1 Experimental Setup
Our evaluation is conducted on the Causal Triplet benchmark (Liu et al., 2023), specifically designed for intervention-centric causal representation learning. This benchmark features three distinct settings of increasing visual complexity: single-object synthetic scenes, multi-object synthetic scenes (both from ProcTHOR (Deitke et al., 2022)), and challenging real-world scenes from Epic-Kitchens (Damen et al., 2022). In all settings models are trained on pairs of pre- and post-intervention images with action labels and are evaluated for their ability to infer the action. Further details on the datasets and data filtering procedures are provided in the Appendix.
@@ 6.1 Experimental Setup
We compare our two proposed models against the baselines from the Causal Triplet paper (Liu et al., 2023), including vanilla ResNets (He et al., 2016), methods incorporating causal regularization (ICM, SMS), and object-centric approaches (Slot Attention (Locatello et al., 2020b), GroupViT (Xu et al., 2022)).
6.2 Main Quantitative Results
Table 1: Single-object ProcTHOR results. Our Global Delta Embedding model significantly improves OOD generalization under both compositional and systematic shifts. (R: ResNet-18, V: ViT-DINO)
@@ 6.2 Main Quantitative Results
[…]
Our CDE framework consistently delivers substantial improvements in OOD accuracy across all evaluation settings, establishing a new state of the art. For single-object scenes, our global CDE model cuts the generalization gap from 0.56 to 0.21 while matching IID accuracy (Table 1). In challenging multi-object and real-world settings (Table 2), our Patch-Wise model outperforms all baselines, including oracle methods that use ground-truth segmentation masks. Notice that the ResNet18 encoder performs significantly worse than the rest of the encoders, possibly indicating that this backbone cannot disentangle latent representations that depend on the actions from the rest of the scene as required by the model of Figure 2.
Table 2: Results across multi-object ProcTHOR and Epic-Kitchens (systematic shift). Oracle-mask utilizes a ground truth mask to isolate the intervened object. Conversely, other approaches must infer the action and object without this supervision.
Dataset Method IID Acc. OOD Acc. Gap
ProcTHOR ResNet $0.83_{\pm 0.01}$ $0.30_{\pm 0.08}$ $0.53$
@@
Ours(ViT-MAE) $0.91_{\pm 0.1}$ $0.30_{\pm 0.02}$ $0.61$
Ours(ViT-DINO) $0.92_{\pm 0.0}$ $0.45_{\pm 0.03}$ $0.47$
[…]
Epic-Kitchens ResNet $0.42_{\pm 0.03}$ $0.17_{\pm 0.03}$ $0.25$
@@
6.3 Action Relationships in Causal Delta Space
To study the semantic structure of the learned delta space, we investigated whether the model could discover fundamental relationships between actions on its own. We computed the pairwise cosine similarity between all learned action representations. The result is visualized in the appendix (Figure 7). The analysis reveals that the model has learned a perfect anti-parallel relationship for opposite actions. The cosine similarity between the representations for open and close, for dirty and clean, as well as for turn on and turn off, is -1.0. This demonstrates that our framework not only separates the action concepts but also discovers opposing relationships between them, organizing the representations in a meaningful way. A similar pattern is observed in the more challenging real-world dataset where the model learns the anti-parallel representations for the open and close action pair, as well as for the fold and stretch pair (see Figure 8 in the Appendix for details).
In summary, the combination of strong predictive properties and consistent semantic structure demonstrates that our CDE framework learns meaningful representations of interventions. For further geometric analysis of the delta space, including UMAP projections, refer to the Appendix.
6.4 Ablation Study
Table 3: Ablation study of our method’s components on the ViT-DINO model. Results are for the single-object systematic shift setting, showing the impact on OOD accuracy when each core component is removed.
@@ 6.4 Ablation Study
   w/o Sparsity Loss $0.96$ $0.73$
   w/o Contrastive Loss $0.95$ $0.68$
   CE Loss only $0.94$ $0.67$
Baseline (Liu et al., 2023) $0.95$ $0.47$
@@
To understand the contribution of each component of our CDE framework, we also conducted a series of ablation studies, by analyzing the impact of each major loss component on the performance of our primary model with a ViT-DINO backbone. Table 3 presents the results, comparing our full model against versions where each loss component is removed, and a baseline trained only with standard CE loss.
@@
The results demonstrate the effectiveness of our approach. Our full model achieves an OOD accuracy of 75.075.0%, a +8+8 point improvement over delta embedding representation trained solely with a CE objective, validating that explicitly structuring the representation space leads generalization improvements. Removing the supervised contrastive loss component causes a $7$-point drop in OOD accuracy. Removing the sparsity loss term causes $2$-point drop. Please refer to the Appendix for further ablation experiments.
7 Conclusion
@@
This paper introduces the Causal Delta Embedding (CDE) framework, a simple yet effective approach to interventional causal representation learning. By explicitly modeling interventions as delta vectors in a structured latent space, CDE inherently satisfies the properties of independence, sparsity and invariance, leading to improved generalization. Our empirical validation on the Causal Triplet challenge demonstrates that CDE achieves state-of-the-art OOD generalization, outperforming prior methods across synthetic and real world datasets. Beyond quantitative gains, we show that CDE learns semantically meaningful representations without supervision, where opposing actions have anti-parallel representations. Despite the promising results, we acknowledge that limitations remain for real-world data, since both IID and OOD accuracies are still low for real world deployment, and also the use of universal delta embeddings for each action limits its ability to capture context-dependent visual transformations of actions. Future research directions includes the development of mechanisms to increase the robustness of the method in real-world scenarios involving increased noise or occlusions, extending the framework to video streams for modeling causal dynamics in temporal sequences, and investigating compositional properties of delta embeddings to enable multi-step interventions and generalization to novel action sequences.
Reproducibility statement
The previous sections have outlined the main building blocks of the proposed method, as well as the approach followed in the experiments, with the Appendix providing additional information and results. The code to reproduce the experiments is attached as supplementary material (without any identifiable information of authors) and will be made publicly available upon acceptance. Finally, all experiments were carried out by strictly following the Causal Triplet benchmark (Liu et al., 2023) evaluation protocols, which relies on the publicly available ProcTHOR (Deitke et al., 2022) and Epic-Kitchens (Damen et al., 2022) datasets.
Acknowledgments
The work leading to these results has been funded by the European Union under Grant Agreement No. 101057821 (project RELEVIUM). Views and opinions expressed are however those of the authors and do not necessarily reflect those of the European Union or the granting authority (HaDEA). Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Dataset Details
The ProcTHOR dataset Deitke et al. (2022) provides synthetic indoor scenes. For our single-object scenes, each scene contains one manipulated object, ensuring a clear focus on the intervention. In multi-object scenes, multiple objects are present, increasing the visual complexity of the scene, although only one object is again manipulated. We follow the dataset generation and filtering procedures as described in Liu et al. (2023) to ensure consistency with the Causal Triplet benchmark.
@@
The Epic-Kitchens dataset Damen et al. (2022) comprises real world egocentric videos of diverse kitchen activities. From this, we extract pre- and post-intervention image pairs. Unlike synthetic environments, Epic-Kitchens introduces significant real world challenges such as camera motion, varying lighting conditions, occlusions and dynamic backgrounds, making the task of isolating interventions particularly challenging. To ensure dataset quality, a two-stage filtering process using Grounding DINO Liu et al. (2024) for zero-shot object detection is applied. For each extracted pair, the pipeline verifies that the target object appears clearly in both frames with a detection confidence above a set threshold $t=0.45$. This automated filtering removes cases with poor object visibility or excessive motion blur.
A.1 Visualizations of OOD Shifts
Figures 4, 5 and 6 visually illustrate the compositional and systematic distribution shifts utilized in the Causal Triplet benchmark.
@@ Appendix B Geometric Analysis of Causal Delta Embeddings
B.1 Action Representation Relationships Learned from Real-World Datasets
Figure 7 illustrates the the pairwise cosine similarities between the embeddings learned for all actions in the ProcTHOR dataset, while Figure 8 presents the same information for the more challenging real-world Epic Kitchens dataset. We observe that in both cases the learned relationships for certain opposing actions such as open and close as well as fold and stretch are antiparallel in the embedding space. Note, however, that the representation does not capture all antiparallel relationships (e.g., add and remove have a similarity of $0.19$).
B.2 Analysis of Learned Action Representations
@@ B.3 UMAP Projection
Figures 9 and 10 present the UMAP projection of individual delta embeddings from the IID and OOD test set of the single-object environment respectively. The delta embeddings in the IID setting achieve a clear separation between each action, leading to a near perfect IID accuracy as was presented by our experiments. On the other hand, while strong intra-class cohesion is visible, the global separation between these action clusters is not always visually distinct in the OOD setting. This suggests that while representations remain locally coherent, action representations are not as clearly discriminated compared to the IID setting. It is worth mentioning, however, that the 2D projection may not fully capture the features of the high-dimensional latent space.
@@ Appendix C Ablation Studies
All the subsequent experiments ran on the single-object systematic shifts setting, except for the Top-K ablation study which ran on the multi-object systematic shifts setting. Also these experiments were performed in a smaller subset of the dataset.
@@
ViT-DINO Ours (CE Only) $0.67$
ViT-DINO Ours (Full Model) 0.75
@@
C.1 Impact of Backbone Architecture
To isolate the contribution of our CDE framework from the choice of feature extractor, we conducted a controlled comparison between our ViT-DINO backbone and the ResNet-18 backbone used in the original Causal Triplet benchmark.
C.2 Impact of Loss Hyperparameters
@@ D.3 Image Augmentations
We do not apply any augmentations to the images, since we do not want to modify the interventional nature of the pairs. Augmentation in this problem could harm our assumptions. For example, a rotation could affect equation 1 and eliminate the faithfulness of the encoder. We leave it as future work to investigate whether augmentations can boost OOD performance under different assumptions. We only resize images to $224\times 224$ pixels and apply zero-mean normalization with unit variance.
D.4 Optimization
@@
We use a batch size of 128 and an AdamW Loshchilov and Hutter (2017) optimizer with a cosine annealing learning scheduler for 50 epochs. In the real world setting, we instead train for 100 epochs. The ViT feature extractor is not frozen but fine-tuned with a reduced learning rate of $10\%$ of the network’s base learning rate, which is set to $1\times 10^{-4}$. The weight decay parameter is $0.05$. All reported results include standard deviations computed over three independent runs with different random seeds.
D.5 Model Architectures
To ensure reproducibility and clarify the pre-training objectives used in our experiments, Table 8 details the specifications for the vision backbones used in our Causal Delta Embedding (CDE) framework and the comparative baselines. We utilize the implementations provided by the timm library.
Table 8: Comparison of Vision Backbones utilized in experiments.
Model Architecture Pre-training Objective Dataset Params
ViT-DINO ViT-S/16 Self-Supervised (Distillation) ImageNet-1k 21.7M
ViT-MAE ViT-B/16 Self-Supervised (Reconstruction) ImageNet-1k 85.8M
CLIP ViT-B/16 Weakly-Supervised (Contrastive) Web-400M $\sim$86M
ResNet-18 ResNet-18 Supervised (Classification) ImageNet-1k 11.7M
Appendix E Actionable counterfactuals and delta embeddings
If we relax the requirement for identical noise across observations ($\epsilon\neq 0$ in equation 3), then the delta embedding of equation 1 becomes (assuming column vectors)
$\delta^{T}_{a}=\begin{bmatrix}\epsilon_{1},&\epsilon_{2},&\dotsc,&\tilde{z}_{a}-z_{a}+\epsilon,&\dotsc,&\epsilon_{l}\end{bmatrix}^{T}$
where $l$ is the number of vector dimensions and $\epsilon_{1}$, $\epsilon_{2}$ etc are assumed to be zero-mean noise variables that do not depend on $a$ or $\tilde{z}_{a}-z_{a}$. We use the following representation for convenience $\delta^{T}_{a}=\begin{bmatrix}\mathbf{u}^{T}_{\epsilon},&\mathbf{u}^{T}_{a}\end{bmatrix}$, where $\mathbf{u}_{\epsilon}$ is the part of the representation that does not depend on $a$ and $\mathbf{u}_{a}$ the part that does. Consider that this representation is used with a binary logistic regression model with parameters $\mathbf{w}^{T}=\begin{bmatrix}\mathbf{w}^{T}_{\epsilon},&\mathbf{w}^{T}_{a}\end{bmatrix}$. The gradient of the binary cross-entropy loss with respect to $\mathbf{w}_{\epsilon}$ is
$\nabla_{\mathbf{w}_{\epsilon}}\left[\mathcal{L}_{CE}\right]=\mathbb{E}\left[\left(\sigma\left(\mathbf{w}^{T}_{\epsilon}\mathbf{u}_{\epsilon}+\mathbf{w}^{T}_{a}\mathbf{u}_{a}+b\right)-a\right)\mathbf{u}_{\epsilon}\right]$
Note that at $\mathbf{w}_{\epsilon}=\mathbf{0}$ this becomes
$\nabla_{\mathbf{w}_{\epsilon}}\left[\mathcal{L}_{CE}\right]|_{\mathbf{w}_{\epsilon}=\mathbf{0}}=\mathbb{E}\left[\left(\sigma\left(\mathbf{w}^{T}_{a}\mathbf{u}_{a}+b\right)-a\right)\mathbf{u}_{\epsilon}\right]$
Given that $\mathbf{u}_{\epsilon}$ is independent of $a$, $\mathbf{u}_{a}$ and $b$ then this equation can be factorized as
$\nabla_{\mathbf{w}_{\epsilon}}\left[\mathcal{L}_{CE}\right]|_{\mathbf{w}_{\epsilon}=\mathbf{0}}=\mathbb{E}_{\mathbf{u}_{a},a}\left[\left(\sigma\left(\mathbf{w}^{T}_{a}\mathbf{u}_{a}+b\right)-a\right)\right]\mathbb{E}_{\mathbf{u}_{\epsilon}}\left[\mathbf{u}_{\epsilon}\right]=\mathbf{0}$
since $\epsilon_{1},\epsilon_{2},\dotsc,\epsilon_{d}$ are zero mean variables. This means that $\mathbf{0}$ is a stationary point of $\mathcal{L}_{CE}$. Since $\mathcal{L}_{CE}$ is convex for logistic regression, for any given $\mathbf{w}_{a}$ and $b$, $\mathbf{w}_{\epsilon}=\mathbf{0}$ is a global minimum of the loss. This implies that under imperfect interventions (actionable counterfactuals) under the zero mean independent noise assumption, the cross-entropy loss is minimized where the nonzero noise variables are ignored. Similar arguments apply in the multiclass case.
Appendix F Adaptation of Object-Centric Models for Action Prediction
To leverage the latent structure of object-centric representations for downstream action reasoning, we adapt the implicit Slot Attention framework (Locatello et al., 2020b). The model decomposes each input scene into a set of $N$ spatially and semantically related regions (slots), each characterized by its own feature vector. Given a pair of pre- and post-intervention images, we explore three distinct aggregation strategies to bridge the object-centric slots with the action encoder.
  • Slot-Avg: This baseline approach performs average-pooling over the $N$ slots for each image independently. The resulting single feature vectors for the pre- and post-intervention images are concatenated and passed to the action encoder, effectively treating the aggregated slots as a distributed representation.
  • Slot-Dense: This strategy densely pairs every slot from the pre-intervention image with every slot from the post-intervention image, resulting in $N\times N$ combinations. All pairs are processed by the action encoder to generate relation embeddings, which are then aggregated via average-pooling to form the final action representation.
  • Slot-Match: This method selectively pairs slots across the two images based on cosine similarity. Only the matched pairs are passed to the action encoder. The resulting $N$ relation embeddings are aggregated using max-pooling to capture the most significant latent changes corresponding to the intervention.